The reaction between potassium chlorate and red phosphorus takes place when you strike a match on a matchbox. If you were to react 51.7 g of potassium chlorate (KClO3) with excess red phosphorus, what mass of tetraphosphorus decaoxide (P4O10) would be produced?
1
Expert's answer
2020-04-18T07:00:39-0400

The reaction is:

10KClO3(s) + 3P4(s) = 3P4O10(s) + 10KCl(s)

Convert to moles:

n(KClO3) = m/M = 51.7g/122.55g/mol = 0.422 mol

Multiply by mole ratio:

n(P4O10) = n(KClO3)*3/10 = 0.127 mol

Convert to grams:

m(P4O10) = n*M = 0.127mol*283.89g/mol = 36.05 g